# Hierarchical Deep Learning for Diatom Image Classification: A Multi-Level Taxonomic Approach

### v1

@@ IV-B Error Robustness and Taxonomic Distance Analysis
TABLE II: Taxonomic Distance Analysis on Species-Level Errors
@@
Table II quantifies error locality through taxonomic distance analysis, defined as the number of edges in the taxonomy tree between the predicted and true species. H-COFGS produces 92.5% of errors at distance 1 (same genus) compared to 67.2% for F-S. The mean taxonomic distance for errors is 1.209 for H-COFGS versus 1.955 for F-S, representing a 38.2% reduction.
@@ V-C Error Locality and Biological Interpretability
Table II shows that H-COFGS concentrates 92.5% of its species-level errors within the same genus (distance = 1), compared to only 67.2% for F-S. The flat baseline produces significantly more errors at distance 4 (same class but different order or family): 22.4% versus 3.0% for H-COFGS. This difference has meaningful implications: hierarchical predictions are more aligned with biological structure, producing mistakes that remain taxonomically local and thus easier for domain experts to interpret and correct.
@@ V-D Limitations
Second, we evaluate only one backbone architecture (EfficientNet-B0). Other architectures such as ResNet, Vision Transformers, or ConvNeXt may exhibit different behavior, and the relative benefits of hierarchical prediction may vary with backbone capacity and inductive biases.
Third, our inference procedure uses greedy hierarchical prediction (sequential argmax). While this guarantees taxonomic consistency, it may propagate errors from upper to lower levels. The appendix explores beam search as an alternative, but more sophisticated inference strategies (e.g., joint decoding with structured prediction) warrant further investigation.
Fourth, the dataset originates from a specific imaging protocol and geographic region. Domain shift—differences in microscope settings, staining procedures, or regional species distributions—remains a challenge for real-world deployment and requires evaluation on external datasets.
@@ V-E Future Directions
Several directions could extend this work. Larger datasets covering more taxonomic groups would enable more robust evaluation of hierarchical architectures. Incorporating semi-supervised or self-supervised learning could leverage abundant unlabeled diatom imagery to improve feature representations.
@@ VI Conclusion
This work examined whether explicitly encoding taxonomic hierarchy in deep learning architectures improves diatom classification compared with flat single-level prediction. We introduced a hierarchical convolutional network with five cascaded heads that jointly predict class, order, family, genus, and species via feature fusion and taxonomy-derived masks that restrict each level to valid descendants. On a filtered dataset of 1,456 images spanning 82 species, the hierarchical model matches flat baselines at the species level while substantially outperforming them at all upper taxonomic levels and producing more taxonomically local errors.
Progressive training experiments suggest that hierarchical masks provide top-down constraints on the prediction space, while losses at all taxonomic levels drive bottom-up gradients through the shared backbone, yielding a hierarchical multi-task objective that improves both accuracy and biological interpretability. These results indicate that hierarchical architectures offer a principled way to embed domain structure into deep learning models for biological monitoring, and motivate future work on larger datasets, richer inference and uncertainty estimation, and applications to other taxa with nested taxonomic structure.
@@ Appendix A Alternative Inference Strategies
The main text uses greedy hierarchical prediction during inference: at each taxonomic level, we select the class with the highest probability (argmax) and use it to construct the mask for the next level. This procedure guarantees taxonomic consistency but may propagate errors from upper to lower levels. Here we evaluate two alternative strategies:
@@
    Level-wise argmax without hierarchical constraints: At each level, select the class with highest probability independently, ignoring taxonomic consistency. This may produce biologically implausible predictions (e.g., a genus not belonging to the predicted family).

### v2

@@ IV-B Error Robustness and Taxonomic Distance Analysis
TABLE III: Taxonomic Distance Analysis on Species-Level Errors
@@
Table III quantifies error locality through taxonomic distance analysis, defined as the number of edges in the taxonomy tree between the predicted and true species. H-COFGS produces 92.5% of errors at distance 1 (same genus) compared to 67.2% for F-S. The mean taxonomic distance for errors is 1.209 for H-COFGS versus 1.955 for F-S, representing a 38.2% reduction.
@@ V-C Error Locality and Biological Interpretability
Table III shows that H-COFGS concentrates 92.5% of its species-level errors within the same genus (distance = 1), compared to only 67.2% for F-S. The flat baseline produces significantly more errors at distance 4 (same class but different order or family): 22.4% versus 3.0% for H-COFGS. This difference has meaningful implications: hierarchical predictions are more aligned with biological structure, producing mistakes that remain taxonomically local and thus easier for domain experts to interpret and correct.
@@ V-D Limitations
Second, we evaluate only one backbone architecture (EfficientNet-B0). Other architectures, such as ResNet, Vision Transformers, or ConvNeXt, may exhibit different behavior, and the relative benefits of hierarchical prediction may vary with backbone capacity and inductive biases.
Third, the dataset originates from a specific imaging protocol and geographic region. Domain shift—differences in microscope settings, staining procedures, or regional species distributions—remains a challenge for real-world deployment and requires evaluation on external datasets.
@@ V-E Future Directions
Several directions could extend this work. Larger datasets covering more taxonomic groups would enable a more robust evaluation of hierarchical architectures. Incorporating semi-supervised or self-supervised learning could leverage abundant unlabeled diatom imagery to improve feature representations.
@@ VI Conclusion
This work examined whether explicitly encoding taxonomic hierarchy in deep learning architectures improves diatom classification compared with flat single-level prediction. We introduced a hierarchical convolutional network with five cascaded heads that jointly predict class, order, family, genus, and species via cascaded probability propagation and taxonomy-derived masks that restrict each level to valid descendants during both training and greedy hierarchical inference. On a filtered dataset of 1,456 images spanning 82 species, the hierarchical model matches flat baselines at the species level while substantially outperforming them at all upper taxonomic levels and producing more taxonomically local errors when fine-grained predictions fail.
Progressive training experiments suggest that hierarchical masks provide top-down constraints on the prediction space, while losses at all taxonomic levels drive bottom-up gradients through the shared backbone, yielding a hierarchical multi-task objective that improves both accuracy and biological interpretability. These results indicate that hierarchical architectures offer a principled way to embed domain structure into deep learning models for diatom classification, and motivate future work on larger datasets, richer inference and uncertainty estimation, and applications to other taxa with nested taxonomic structure.
@@ Appendix A Alternative Inference Strategies
The main text uses greedy hierarchical prediction during inference: at each taxonomic level, we select the category with the highest probability (argmax) and use it to construct the mask for the next level. This procedure guarantees taxonomic consistency but may propagate errors from upper to lower levels. Here we evaluate two alternative strategies:
@@
    Level-wise argmax without hierarchical constraints: At each level, select the category with the highest probability independently, ignoring taxonomic consistency. This may produce biologically implausible predictions (e.g., a genus not belonging to the predicted family).